OpenStudy (anonymous):

A computer can sort x objects in t seconds, as modeled by the function below: How many objects are required to keep the computer busy for exactly 9 seconds? Round to the nearest whole object.

OpenStudy (isaiah.feynman):

Where is the function?

OpenStudy (dape):

Put in t=9 and solve for x using the quadratic formula.

OpenStudy (anonymous):

that gives me x=0 x=−0.42857142857142855 right?

OpenStudy (dape):

I'm not getting those answers, could you show me your steps so I can help you clear up any confusion =)

OpenStudy (anonymous):

so if i put in \[ x=\frac{ -b+/-\ \sqrt{b^2-4ac} }{ 2a }\] and I put in a=.007 b=003 c=0 the two answers I get are x=0 x=−0.42857142857142855

OpenStudy (dape):

Ah, I see, the c should be -9 in this case, since we start with \[9=0.007x^2+0.003x\] which is the same as \[0.007x^2+0.003x-9=0\]

OpenStudy (anonymous):

so the answer shoud be 36

OpenStudy (dape):

Yes, you're right!